# Space-time Peer-to-Peer Distribution of Multi-party Entanglement for Any Quantum Network

### v1

@@ 5. DISTRIBUTING ANY QUANTUM COMPUTATION
(13) $|\psi_{s}\rangle=\alpha|0\rangle+\beta|1\rangle,\,\alpha\beta^{*}\in i\mathbbm{R}$
@@
Now, go back to the special case where we want to share a quantum secret among $N$ remote clients. If one thinks of the circuits where the secret is always being split at the center node as shown in Fig. 10(a), one notices that all the Hadamard gates between the CZ gates are canceled with each other, and thus all CZ gates form a layer that corresponds to a star graph, which justifies our statement earlier. The distribution of a star graph with the (ST-)P2PGSD algorithm also has exactly the same pattern as the cascading [2,2] sharing scheme in the sense that the required bell pairs span a tree that covers all client nodes, which offers an alternative understanding of our previously described P2P quantum secret sharing strategy. We remark that the star graph has also been proven to be the fundamental resource for the quantum sharing of classical secrets as well as the splitting of quantum information (Karlsson et al., 1999; Hillery et al., 1999), which can also be viewed as a special case of our distributed quantum computation strategy. However, the P2P quantum secret sharing scheme presented here is different from these protocols in the sense that no information can be retrieved for the quantum state if only a proper subset of the clients cooperates together.
@@ Acknowledgements.
We thank Antonio Barbalace and Allen Zang for their suggestions and acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209, FA9550-23-1-0338), DARPA (HR0011-24-9-0359, HR0011-24-9-0361), NSF (OMA-1936118, ERC-1941583, OMA-2137642, OSI-2326767, CCF-2312755), NTT Research, Packard Foundation (2020-71479), and the Marshall and Arlene Bennett Family Research Program. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and Advanced Scientific Computing Research (ASCR) program under contract number DE-AC02-06CH11357 as part of the InterQnet quantum networking project. This work was also completed with resources provided by the University of Chicago’s Research Computing Center and part of this research was performed while the author was visiting the Institute for Mathematical and Statistical Innovation (IMSI), which is supported by the National Science Foundation (Grant No. DMS-1929348).
@@ Appendix A PROOF OF THE THEOREMS AND STATEMENTS
It is easy to see that MGST guarantees to solve $\mathscr{P}$ within $|V_{S}|$ shots. Let $[\mathscr{b}_{|V_{S}|}]$ denote its flow solution. It suffices to prove that we can modify it into another solution $[\mathscr{b}^{\prime}_{|V_{S}|}]$ with no channel containing flow greater than $\lfloor\frac{V_{S}}{2}\rfloor$. Find any channel that contains flow greater than $\lfloor\frac{|V_{S}|}{2}\rfloor$, then use the flow-decomposition theorem to get all the directed paths. Since the paths that utilize the same channel will have the same direction as the theorem guarantees, one can easily show that removing or reversing a subset of paths passing a specific node will not increase the absolute flow values in the channels connecting to it. There are $f_{1}>\lfloor\frac{|V_{S}|}{2}\rfloor$ paths passing this channel to more than $\lfloor\frac{|V_{S}|}{2}\rfloor$ targets from the roots. All other $f_{2}=|V_{S}|-f_{1}\leq\lfloor\frac{|V_{S}|}{2}\rfloor$ paths do not use such channel. Then move the root to the downstream endpoint of this channel. Now, we just need to reverse the direction of any $f_{2}$ paths within $f_{1}$ paths and remove other $f_{1}-f_{2}$ paths to get the new flow solution. Such operation will not create any new channels that carry flow greater than $\lfloor\frac{|V_{S}|}{2}\rfloor$ as argued above while making the selected channel now consume only $f_{2}$ flow. This process can be repeated until no channels contain flow greater than $\frac{|V_{S}|}{2}$. And one can easily construct an example $\mathscr{P}$ that requires at least $\frac{|V_{S}|}{2}$ shots to distribute by assuming the graph state consists of merely bell pairs and the network has a single bottleneck channel with width $1$, which proves that this bound is tight and optimal. We remark that this result is the first tight upper-bound, to our best knowledge.
@@ Appendix B RECOVERY ALGORITHM
As for recovering using memory, a naive and simple idea is just always to store the qubits as long as there is enough long-term memory. This will indeed lower the shot usage to establish a certain path; however, it will not only cause a blow-up in the usage of cumulative memory but may also waste the resources if a high probability link is stored. One observes that there is certainly some threshold behavior in terms of cumulative memory usage, and we shall illustrate it via an intuitive but analytically solvable example here: the task is to establish a bell pair link along a path with $n$ channels, which has identical probability $p$. We consider a simple storing strategy that treats two shots as one cycle and always saves one qubit if the two switched qubits for each node at the first shot and the qubits at two ends must be saved for two shots. We can calculate the cumulative memory usage to execute such a cycle as #$\#\mathscr{C}_{c2}=n+3$, while the expected number of cycles to be executed can be approximated as $T_{c2}\approx\frac{1}{p^{n}(2-p)^{n}}$, from which we calculate the expected cumulative memory usage as #$\#\mathscr{C}_{\mathrm{tot},\,c2}=\frac{n+3}{p^{n}(2-p)^{n}}$. Compared to the no-saving strategy, whose expected cumulative memory usage is $\frac{2}{p^{n}}$, one immediately deduces a threshold probability as $p_{\mathrm{th}}=2-\left(\frac{n+3}{2}\right)^{\frac{1}{n}}$, in case $n=2$, this corresponds to $p_{th}\approx 0.42$, above which one concludes that the no-saving strategy has an advantage. Considering a cycle with $k>2$ shots with an adaptive saving strategy is much more tricky since the neighboring links can be merged into one link to avoid saving the intermediate qubits, although one can easily calculate the expected number of cycles as $T_{ck}\approx\frac{1}{(1-(1-p)^{k})^{n}}$. Even worse, if the memory link is probabilistic in a more general case, the analytical calculation of the expected number of cycles also becomes hardly possible since one can show that this corresponds to a percolation problem in a grid network (Pant et al., 2019).

### v3

@@ 5. DISTRIBUTING ANY QUANTUM COMPUTATION
(13) $|\psi_{s}\rangle=\alpha|0\rangle+\beta|1\rangle,\,\alpha\beta^{*}\in i\mathbb{R}$
@@
Now, go back to the special case where we want to share a quantum secret among $N$ remote clients. If one thinks of the circuits where the secret is always being split at the center node as shown in Fig. 10(a), one notices that all the Hadamard gates between the CZ gates are canceled with each other, and thus all CZ gates form a layer that corresponds to a star graph, which justifies our statement earlier. The distribution of a star graph with the (ST-)P2PGSD algorithm also has the exact same pattern as the cascading [2,2] sharing scheme in the sense that the required Bell pairs span a tree that covers all client nodes, which offers an alternative understanding of our previously described P2P quantum secret sharing strategy. We remark that the star graph has also been proven to be the fundamental resource for the quantum sharing of classical secrets as well as the splitting of quantum information (Karlsson et al., 1999; Hillery et al., 1999), which can also be viewed as a special case of our distributed quantum computation strategy. However, the P2P quantum secret sharing scheme presented here is different from these protocols in the sense that no information can be retrieved for the quantum state if only a proper subset of the clients cooperates together.
@@ Acknowledgements.
We thank Chen Qian, Antonio Barbalace and Allen Zang for their suggestions and acknowledge support from the ARO(W911NF-23-1-0077), ARO MURI (W911NF-21-1-0325), AFOSR MURI (FA9550-19-1-0399, FA9550-21-1-0209, FA9550-23-1-0338), DARPA (HR0011-24-9-0359, HR0011-24-9-0361), NSF (OMA-1936118, ERC-1941583, OMA-2137642, OSI-2326767, CCF-2312755), NTT Research, Packard Foundation (2020-71479), and the Marshall and Arlene Bennett Family Research Program. This material is based upon work supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and Advanced Scientific Computing Research (ASCR) program under contract number DE-AC02-06CH11357 as part of the InterQnet quantum networking project. This work was also completed with resources provided by the University of Chicago’s Research Computing Center and part of this research was performed while the author was visiting the Institute for Mathematical and Statistical Innovation (IMSI), which is supported by the National Science Foundation (Grant No. DMS-1929348).
@@ Appendix A PROOF OF THE THEOREMS AND STATEMENTS
It is easy to see that MGST guarantees to solve $\mathscr{P}$ within $|V_{S}|$ shots. Let $[\mathscr{b}_{|V_{S}|}]$ denote its flow solution. It suffices to prove that we can modify it into another solution $[\mathscr{b}^{\prime}_{|V_{S}|}]$ with no channel containing flow greater than $\lfloor\frac{V_{S}}{2}\rfloor$. Find any channel that contains flow greater than $\lfloor\frac{|V_{S}|}{2}\rfloor$, then use the flow-decomposition theorem to get all the directed paths. Since the paths that utilize the same channel will have the same direction as the theorem guarantees, one can easily show that removing or reversing a subset of paths passing a specific node will not increase the absolute flow values in the channels connecting to it. There are $f_{1}>\lfloor\frac{|V_{S}|}{2}\rfloor$ paths passing this channel to more than $\lfloor\frac{|V_{S}|}{2}\rfloor$ targets from the roots. All other $f_{2}=|V_{S}|-f_{1}\leq\lfloor\frac{|V_{S}|}{2}\rfloor$ paths do not use such channel. Then move the root to the downstream endpoint of this channel. Now, we just need to reverse the direction of any $f_{2}$ paths within $f_{1}$ paths and remove other $f_{1}-f_{2}$ paths to get the new flow solution. Such operation will not create any new channels that carry flow greater than $\lfloor\frac{|V_{S}|}{2}\rfloor$ as argued above while making the selected channel now consume only $f_{2}$ flow. This process can be repeated until no channels contain flow greater than $\frac{|V_{S}|}{2}$. And one can easily construct an example $\mathscr{P}$ that requires at least $\frac{|V_{S}|}{2}$ shots to distribute by assuming the graph state consists of merely Bell pairs and the network has a single bottleneck channel with width $1$, which proves that this bound is tight and optimal. We remark that this result is the first tight upper bound, to our best knowledge.
@@ Appendix B RECOVERY ALGORITHM
As for recovering using memory, a naive and simple idea is just always to store the qubits as long as there is enough long-term memory. This will indeed lower the shot usage to establish a certain path; however, it will not only cause a blow-up in the usage of cumulative memory but may also waste resources if a high probability link is stored. One observes that there is certainly some threshold behavior in terms of cumulative memory usage, and we shall illustrate it via an intuitive but analytically solvable example here: the task is to establish a Bell pair link along a path with $n$ channels, which has identical probability $p$. We consider a simple storing strategy that treats two shots as one cycle and always saves one qubit if the two switched qubits for each node at the first shot and the qubits at two ends must be saved for two shots. We can calculate the cumulative memory usage to execute such a cycle as #$\#\mathscr{C}_{c2}=n+3$, while the expected number of cycles to be executed can be approximated as $T_{c2}\approx\frac{1}{p^{n}(2-p)^{n}}$, from which we calculate the expected cumulative memory usage as #$\#\mathscr{C}_{\mathrm{tot},\,c2}=\frac{n+3}{p^{n}(2-p)^{n}}$. Compared to the no-saving strategy, whose expected cumulative memory usage is $\frac{2}{p^{n}}$, one immediately deduces a threshold probability as $p_{\mathrm{th}}=2-\left(\frac{n+3}{2}\right)^{\frac{1}{n}}$, in case $n=2$, this corresponds to $p_{th}\approx 0.42$, above which one concludes that the no-saving strategy has an advantage. Considering a cycle with $k>2$ shots with an adaptive saving strategy is much more tricky since the neighboring links can be merged into one link to avoid saving the intermediate qubits, although one can easily calculate the expected number of cycles as $T_{ck}\approx\frac{1}{(1-(1-p)^{k})^{n}}$. Even worse, if the memory link is probabilistic in a more general case, the analytical calculation of the expected number of cycles also becomes hardly possible since one can show that this corresponds to a percolation problem in a grid network (Pant et al., 2019).